OpenStudy (jenniferjuice):

Solve the equation 1 over 2 x + 2 = 22 for x.

OpenStudy (ivancsc1996):

\[\frac{ 1 }{ 2x+2 }=22 \rightarrow \frac{ \frac{ 1 }{ 22 }-2 }{ 2 }=x=-\frac{ 43 }{ 44 }\]

OpenStudy (anonymous):

flip it and get \[2x+2=\frac{1}{22}\] then solve in two steps

OpenStudy (jenniferjuice):